Let T : R3 → R3 be defined by T (x1

, x2

, x3

) = (x1

, x2

,−x1 − x2

). Find a

matrix which represents T


1
Expert's answer
2021-06-28T16:37:12-0400

Let "T:R^3\to R^3" be defined by "T(x_1, x_2, x_3)=(x_1, x_2, -x_1-x_2)"



"T\begin{pmatrix}
 \begin{bmatrix}
 x_1\\
 x_2 \\
x_3
\end{bmatrix} 
\end{pmatrix}=\begin{bmatrix}
 x_1 \\
 x_2\\
-x_1-x_2
\end{bmatrix}"

"T\begin{pmatrix}
 \begin{bmatrix}
 1\\
 0 \\
0
\end{bmatrix} 
\end{pmatrix}=\begin{bmatrix}
 1 \\
 0\\
-1
\end{bmatrix}"

"T\begin{pmatrix}
 \begin{bmatrix}
 0\\
 1 \\
0
\end{bmatrix} 
\end{pmatrix}=\begin{bmatrix}
 0 \\
 1\\
-1
\end{bmatrix}"


"T\begin{pmatrix}
 \begin{bmatrix}
 0\\
 0 \\
1
\end{bmatrix} 
\end{pmatrix}=\begin{bmatrix}
 0\\
 0\\
0
\end{bmatrix}"

Using these as our columns, the standard matrix for "T" is:


"A=\begin{bmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 -1 & -1 & 0 \\
\end{bmatrix}"